OpenStudy (anonymous):

I have the this function: y=x^(1/2) and I need to calculate the average rate of change. can someon help me?

OpenStudy (anonymous):

I got \[\frac{ 1 }{ \Delta x }\] but the answer in book is \[\frac{ \Delta y }{ \Delta x }=\frac{ 1 }{ \sqrt{x0} + \sqrt{x0+ \Delta x}}\]

OpenStudy (phi):

by average rate of change they mean the slope \( \frac{\Delta y}{\Delta x}\) if you pick a point \(x_0 \) its \(y_0\) value will be \(\sqrt{x_0}\) if you go a small distance \(\Delta x \) so that \(x_1=x_0 +\Delta x\) and \(y_1= \sqrt{x_1}= \sqrt{x_0+\Delta x} \) you can write the average rate of changes as \[ \frac{y_1 -y_0}{x_1- x_0} = \frac{\sqrt{x_0+\Delta x} - \sqrt{x_0}}{\Delta x}\] so far, this looks ugly, because we do not want a \(\Delta x \) in the denominator (we want to let \(\Delta x \) approach 0) however, if you multiply top and bottom by the conjugate (note the plus sign) \[ \sqrt{x_0+\Delta x} + \sqrt{x_0} \] you get a nice result.

OpenStudy (anonymous):

Thank you Phi. I have to study too much both, english and mathematics. 8) thank you very much.

OpenStudy (anonymous):

I understood the solution but how did you get \[\sqrt{x0+\Delta x} + \sqrt{x0}\] ?

OpenStudy (phi):

\[ \frac{y_1 -y_0}{x_1- x_0} = \frac{\sqrt{x_0+\Delta x} - \sqrt{x_0}}{\Delta x} \cdot \frac{\sqrt{x_0+\Delta x} + \sqrt{x_0}}{\sqrt{x_0+\Delta x} + \sqrt{x_0}}\] \[= \frac{x_0+\Delta x - x_0}{\Delta x \left(\sqrt{x_0+\Delta x} + \sqrt{x_0}\right)} \] \[ = \frac{1}{ \sqrt{x_0+\Delta x} + \sqrt{x_0}} \] I don't understand your question **but how did you get \( \sqrt{x0+\Delta x} + \sqrt{x0}\) ? If you are asking why do we multiply top and bottom by that expression, it's because multiplying by a conjugate is a "known trick" (to those who know it) to get rid of the radicals

OpenStudy (anonymous):

that was my question. You need to get some experience ( and many exercices) to find this solutions. At this point, this site is (almost) perfect, because here we can get help when we need it. Where else I would get help in mathematics on sunday? 8) Phi, Thank you very much for your help. I hope you can help me many more times. []s